OpenStudy (anonymous):

solve. -1/5y=10 A.50 B.-50 C.100 D.-100

OpenStudy (anonymous):

-1/5y = 10 multiply the equation by 5 -1y = 50 or just -y = 50 now multiply by -1 to make y positive y = -50 check... -1/5(-50) = 10 -- negative times negative = positive 50/5 = 10 10 = 10 (correct) y = -50

OpenStudy (anonymous):

little tip...when you have a fraction in the equation, if you multiply the equation by the common denominator, it will get rid of your fractions and turn them to whole numbers.

OpenStudy (austinl):

Not to interrupt, but when you have a fraction with a variable, and you need to divide it out. You can just multiply both sides by it's reciprocal. In this case it happens to be 5 :) Carry on :)

OpenStudy (anonymous):

nicely said austin :)

OpenStudy (austinl):

Lets say for example we have, \(\displaystyle \frac{3}{4}x+4=10\) We can subtract 4 from both sides, \(\displaystyle\frac{3}{4}x=6\) Then we can multiply by the reciprocal of 3/4 which is 4/3. \(\displaystyle x=6\cdot\frac{4}{3}=\frac{6\times4}{3}=\frac{24}{3}=8\) Make sense?

OpenStudy (anonymous):

I do it a little different.... 3/4x + 4 = 10 (I would multiply this by 4) 3x + 16 = 40 3x = 40 - 16 3x = 24 x = 8

OpenStudy (austinl):

That is one of the many reasons why I absolutely love algebra, there are so many different routes you can take on your way to the correct answer! :D

OpenStudy (anonymous):

thats just because I do not like dealing with fractions

OpenStudy (austinl):

Fractions are your friend :P My teachers have always said that :)

OpenStudy (anonymous):

Fractions are your friends....I don't know about that...I prefer whole numbers